OpenStudy (anonymous):

PLEASE HELP Use the quadratic formula to solve 2y^2 – 10y + 8 = 0. A. {–4, –1} B. {4, –1} C. {–4, 1} D. {4, 1}

OpenStudy (anonymous):

@GotToLoveJenny

OpenStudy (anonymous):

For an expression of the form: ax² + bx + c = 0, the quadratic formula is: x = [ -b ± √(b² - 4ac) ] / 2a In your case, the variable is y, but it is the same thing. 2y² - 10y + 8 = 0 a = 2 b = -10 c = 8 y = [ -b ± √(b² - 4ac) ] / 2a y = [ -(-10) ± √((-10)² - 4(2)(8)) ] / 2(2) y = [ 10 ± √(100 - 64) ] / 4 y = [ 10 ± √36 ] / 4 y = [ 10 ± 6 ] /4 Therefore, two cases: y = (10 - 6) / 4 = 4/4 = 1 or y = (10 + 6) / 4 = 16/4 = 4 Double-check: If you factor that, you get: 2(y - 1)(y - 4) = 0 and the same answers for y = 1 or y = 4 Also, if you plug in the values, you should get a true statement: y = 1: 2(1)² - 10(1) + 8 =? 0 2 - 10 + 8 = 0 (correct) y = 4: 2(4)² - 10(4) + 8 =? 0 32 - 40 + 8 = 0 (correct) Answer: y = 1 or y = 4

OpenStudy (anonymous):

Use the quadratic formula to solve 2y2 – 3y = 2. A. {–½, –2} B. {–½, 2} C. {½, –2} D. {½, 2}

OpenStudy (anonymous):

do you need help on the first on or second one?

OpenStudy (anonymous):

second

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

y={-1/2, 2} Correct Answer B.